OpenStudy (anonymous):

Solve for y explicitly: dy/dx = 2y/x - (x^2)(y^2) Show the most crucial steps please.

OpenStudy (anonymous):

this is a nonlinear difference equation right?

OpenStudy (anonymous):

do you have any boundary conditions that way you can use taylor series expansion

OpenStudy (anonymous):

It's a Bernoulli equation which you convert to a linear equation and then solve it.

OpenStudy (anonymous):

hmm for final answer in terms of y and x \[\frac{1}{3}* \ln (\frac{y^3 - x^3}{x^3}) -lnx +C=0\]

OpenStudy (anonymous):

No, the correct answer is: y = 2 / (Cx - x^3) You treated the equation as a homogenous one, it doesnt look like you can do that. I solved it by bernoulli and linear solution.

OpenStudy (anonymous):

Okay hmm my bad idk about bernoulli method though

OpenStudy (anonymous):

Bernoulli are differentialequations on the form: dy/dx + P(x)y = Q(x)y^n Where n > 1or else it's just a linear equation. You solve it by dividing the equation by y^n, substituting v = y^1-n, and then solve it linearly.